Walter is a waiter at the Towne Diner. He earns a daily wage of $50, plus tips that are equal to 15% of the total cost of the dinners he serves. What was the total cost of the dinneres he served if he earned $170 on Tuesday?
Answer by mananth(16946)

D(w) = 50 + 0.15 x
where x is the cost of dinner he serves
he earned $170
170=50+0.15x
170-50 = 0.15x
120=0.15x
120/0.15=x
800=x
$800 dinner was served by the waiter.
m.ananth@hotmail.ca